A gas sample with a volume of 525 mL and temperature of -25.0 °C is heated to 175.5 °C.

What is the new volume of the gas if pressure and number of moles are held constant?


1
Expert's answer
2021-10-14T04:03:41-0400

"V_1 = 525 \;mL \\

T_1= -25.0 \;°C = 248 \;K \\

T_2 = 175.5 \;°C = 448.5 \;K \\

P= constant \\

\frac{V_1}{T_1} = \frac{V_2}{T_2} \\

V_2 = \frac{V_1 \times T_2}{T_1} \\

= \frac{525 \times 448.5}{248} = 949.44 \;mL"

Answer: 949.44 mL